OpenStudy (anonymous):

Write 8√19 as an entire radical? Please help...

OpenStudy (lgbasallote):

\[8\sqrt{19}\] wrtite as an entire radical?

OpenStudy (anonymous):

yea please...

OpenStudy (lgbasallote):

what's an entire radical?

OpenStudy (lgbasallote):

oh put 8 inside the radical huh

OpenStudy (anonymous):

\[8 \times \sqrt{19}\] You have to find out: the square root of what number is 8?

OpenStudy (anonymous):

So reverse the operation and you get 64, and you can stick the 64 until a radical symbol

OpenStudy (anonymous):

under*